# Tradeoffs of Diagonal Fisher Information Matrix Estimators

### v2

@@ 3 Variance of Diagonal FIM Estimators
${\mathcal{V}}_{j}(\theta_{i}\,|\,\bm{x})\vcentcolon=\mathrm{Var}\left(\hat{\mathcal{I}}_{j}(\theta_{i}\,|\,\bm{x})\right),$
@@
where $\mathrm{Var}(\cdot)$ means the variance and ^$\hat{\mathcal{I}}_{j}(\theta_{i}\,|\,\bm{x})$ ($j\in\{1,2\}$) is defined by Eqs. 5 and 6 with $\bm{x}=\bm{x}_{1}=\cdots=\bm{x}_{N}$. An estimator with a smaller variance indicates that it is more accurate and more likely to be close to the true FIM. Based on the variance, one can derive sample complexity bounds of the diagonal FIM via Chebyshev’s inequality, see for instance Soen & Sun (2021, Section 3.4).
@@
By its definition, ${\mathcal{V}}_{j}(\theta_{i}\,|\,\bm{x})$ has a simple closed form Soen & Sun (2021), which is detailed in Appendix A. Given a fixed $\bm{x}\in\Re^{I}$, both ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ and ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ have an order of $\mathcal{O}(1/N)$, with $N$ denoting the number of samples of $\bm{y}_{k}$. They further depend on two factors: ① the derivatives of the parameter-output mapping $\bm{\theta}\to\bm{h}$ stored in a $T\times\dim(\bm{\theta})$ matrix, either $\partial_{i}\bm{h}^{a}(\bm{x})$ or $\partial^{2}_{ii}{\bm{h}}^{a}(\bm{x})$, where the latter can be expensive to calculate; and ② the central moments of $\bm{t}(\bm{y})$, whose computation only scales with $T$ (the number of output units) and is independent to $\dim(\bm{\theta})$.
@@
From an information geometry Amari (2016) perspective, $\mathcal{I}(\bm{\theta})$, ${\mathcal{V}}_{1}(\bm{\theta})$, and ${\mathcal{V}}_{2}(\bm{\theta})$ are all pullback tensors of different orders. For example, $\mathcal{I}(\bm{\theta})$ is the pullback tensor of $\mathcal{I}(\bm{h})$ and the singular semi-Riemannian metric Sun & Nielsen (2019). They induce the geometric structures of the neuromanifold (parameterized by $\bm{\theta}$) based on the corresponding low dimensional structures of the exponential family (parameterized by $\bm{h}$).
@@
Table 1: Exponential family statistics with eigenvalue upper bounds for central moments. For the classification case, $\sigma(\bm{x})$ denotes the softmax value of the logit $\bm{h}(\bm{x})$. † denotes exact eigenvalues rather than upper bounds.
@@
Learning Setting Exponential Family Output $\mathcal{Y}$ Sufficient Statistic $\bm{t}(\bm{y})$ UB $\lambda_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))$ UB $\tilde{\lambda}_{\max}(\mathcal{K}(\bm{t}\,|\,\bm{x}))$
Regression (Isotropic) Gaussian $\Re^{T}$ ${\bm{y}}$ $1$ $3$
@@ 4 Practical Variance Estimation
To further understand the dependencies of the derivative and central moment terms, the FIM and the variance of each estimator can be bounded to strengthen intuition. First, we present simple bounds for the FIM $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ and variances of estimators ^$\hat{\mathcal{I}}_{j}(\theta_{i}\,|\,\bm{x})$.
@@
$\frac{1}{N}\cdot\|\partial^{2}_{ii}{\bm{h}}(\bm{x})\|^{2}_{2}\cdot\lambda_{\min}(\mathcal{I}(\bm{h}\,|\,\bm{x}))$ $\leq{\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ $\leq\frac{1}{N}\cdot\|\partial^{2}_{ii}{\bm{h}}(\bm{x})\|^{2}_{2}\cdot\lambda_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x})),$ (12)
where $\mathcal{M}=\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})-\mathcal{I}({\bm{h}}\,|\,\bm{x})\otimes\mathcal{I}({\bm{h}}\,|\,\bm{x})$; $\lambda_{\min}(\cdot)$ / $\lambda_{\max}(\cdot)$ denotes the minimum / maximum matrix eigenvalue; and given a 4-dimensional tensor $\mathcal{T}\in\Re^{T\times T\times T\times T}$, $\tilde{\lambda}_{\min}(\mathcal{T})$ and $\tilde{\lambda}_{\max}(\mathcal{T})$ are solutions to the following variational equations:
[…]
To help ground Theorem 4.1, in Table 1, we summarize different sufficient statistics quantities for common learning settings — with further learning setting implications presented in Section 5. We further note that Eqs. 11 and 12 (and many subsequent results) can be further generalized for the variance of off-diagonal elements, see Appendix B. In this generalization, the exact form of the eigenvalue / spectral quantities change, but the dependence on the neural network weights split according the FIM coordinate, i.e., in Eq. 11 $\|\partial_{i}{\bm{h}}(\bm{x})\|^{4}_{2}$ becomes $\|\partial_{i}{\bm{h}}(\bm{x})\|^{2}_{2}\cdot\|\partial_{j}{\bm{h}}(\bm{x})\|^{2}_{2}$ and in Eq. 11 $\|\partial^{2}_{ii}{\bm{h}}(\bm{x})\|^{2}_{2}$ becomes $\|\partial^{2}_{ij}{\bm{h}}(\bm{x})\|^{2}_{2}$.
We stress that the values $\tilde{\lambda}_{\min}(\mathcal{T})$ and $\tilde{\lambda}_{\max}(\mathcal{T})$ in Eq. 13 correspond to tensor eigenvalues if and only if $\mathcal{T}$ is a supersymmetric tensor Lim (2005) (aka totally symmetric tensor), i.e., if its indices are permutation invariant. In this case, Eq. 13 is exactly the definition of the maximum and minimum Z-eigenvalues. The variational form mirrors the Courant-Fischer min-max theorem for finding the eigenvalues of symmetric matrices Tao (2012). In the case of Eq. 11, with $\mathcal{M}=\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})-\mathcal{I}({\bm{h}}\,|\,\bm{x})\otimes\mathcal{I}({\bm{h}}\,|\,\bm{x})$, the tensor is not a supersymmetric tensor in general. Despite this, we note that the lower bound of Eq. 11 is non-trivial.
Nevertheless, a weaker bound as compared to Eq. 11 can be established based on the tensor Z-eigenvalue.
@@
The tensor eigenvalue is typically expensive to calculate. However in our case, the eigenvalues $\tilde{\lambda}_{\min}(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x}))$ and $\tilde{\lambda}_{\max}(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x}))$ on the RHS of Eqs. 14 and 15 can be calculated via Kolda & Mayo (2014)’s method with $\mathcal{O}(T^{4}/4!)$ complexity. In this paper, we assume $T$ is reasonably bounded and are mainly concerned with the complexity w.r.t. $\dim(\bm{\theta})$. From this perspective, all our bounds scale linearly w.r.t. $\dim(\bm{\theta})$, and thus can be computed efficiently.
When $\bm{t}(y)-\bm{\eta}(\bm{x})$ is bounded (e.g. in classification), we can upper bound $\tilde{\lambda}_{\max}\left(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})\right)$ with $\lambda_{\max}\left(\mathcal{I}({\bm{h}}\,|\,\bm{x})\right)$, which is easier to calculate.
@@
Suppose $\|\bm{t}(\bm{y})-\bm{\eta}(\bm{x})\|^{2}_{2}\leq B$. Then,
$\tilde{\lambda}_{\max}\left(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})\right)\leq B\cdot\lambda_{\max}\left(\mathcal{I}({\bm{h}}\,|\,\bm{x})\right)\leq B^{2}.$ (16)
It should be noted that as long as the sufficient statistics is bounded $\|\bm{t}\|_{2}$, we have that $B=4\cdot\|\bm{t}\|_{2}^{2}<\infty$. A similar lower bound (without bounded assumption) can be established for the minimum tensor eigenvalue $\tilde{\lambda}_{\min}\left(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})\right)\geq\lambda^{2}_{\min}\left(\mathcal{I}({\bm{h}}\,|\,\bm{x})\right)$, but this ends up being trivial when applying Corollary 4.2’s lower bound, Eq. 14.
One can observe a number of trade-offs by examining Theorem 4.1. An immediate observation is that the first order gradients of $\bm{h}(\bm{x})$ correspond to the robustness of $\bm{h}$ to parameter misspecification (w.r.t. an input $\bm{x}$). As such, from the bounds of Eqs. 10 and 12, the scale of $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ and ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ will be large when small shifts in parameter space yield large changes in the output $\bm{h}(\bm{x})$. Another observation is how the scale of the minimum and maximum eigenvalues of $\mathcal{I}(\bm{h}\,|\,\bm{x})$ affects the scale of $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ and the estimator variances. In particular, when $\lambda_{\min}\left(\mathcal{I}(\bm{h}\,|\,\bm{x})\right)$ increases, the scale of ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ decreases but the scale of $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ and ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ increases. When $\lambda_{\max}\left(\mathcal{I}(\bm{h}\,|\,\bm{x})\right)$ decreases, then the opposite scaling occurs. With these two observations, there is a tension in how the scale of $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ follows the different variances ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ and ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$. The element-wise FIM follows ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ in terms of the scale of neural network derivatives $\|\partial_{i}\bm{h}(\bm{x})\|_{2}$; however, the element-wise FIM follows ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ in terms of the spectrum of sufficient statistics moment $\mathcal{I}(\bm{h}\,|\,\bm{x})$.

### v3

@@ 3 Variance of Diagonal FIM Estimators
^$\hat{\mathcal{I}}_{2}(\theta_{i})$ =.(^$\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\left(\hat{\mathcal{I}}_{2}(\bm{\theta})\right)_{ii}=\frac{1}{N}\sum_{k=1}^{N}\left(\frac{\partial^{2}F(\bm{h}(\bm{x}_{k}))}{\partial^{2}\theta_{i}}-\frac{\partial^{2}\bm{h}^{a}(\bm{x}_{k})}{\partial^{2}\theta_{i}}\cdot\bm{t}_{a}(\bm{y}_{k})\right).$
Correspondingly, the $i$’th diagonal entry of the FIM $\mathcal{I}(\bm{\theta})$, which is the expected value of ^$\hat{\mathcal{I}}_{1}(\theta_{i})$ and ^$\hat{\mathcal{I}}_{2}(\theta_{i})$, is denoted as $\mathcal{I}(\theta_{i})$. Notation is abused in $\mathcal{I}(\theta_{i})$, ^$\hat{\mathcal{I}}_{1}(\theta_{i})$, and ^$\hat{\mathcal{I}}_{2}(\theta_{i})$ as they depend on the whole $\bm{\theta}$ vector rather than solely on $\theta_{i}$. Clearly ^$\hat{\mathcal{I}}_{1}(\theta_{i})\geq 0$, while there is no guarantee for ^$\hat{\mathcal{I}}_{2}(\theta_{i})$ which can be negative. Our results will be expressed in terms of the (central) moments of $\bm{t}(\bm{y})$:
$\bm{\eta}_{a}(\bm{x})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x})}[\bm{t}_{a}(\bm{y})];\quad\quad\quad\mathcal{I}(\bm{h}\,|\,\bm{x})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x})}[(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))^{\top}];$
$\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x})}\left[(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))\otimes(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))\otimes\;(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))\otimes(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))\right],$
where $\otimes$” denotes the tensor product. We denote the covariance of $\bm{t}$ w.r.t. to $p(\bm{y}\,|\,\bm{x})$ as $\mathrm{Cov}^{p}(\bm{t}\,|\,\bm{x})$ — noting that $\mathcal{I}(\bm{h}\,|\,\bm{x})=\mathrm{Cov}^{p}(\bm{t}\,|\,\bm{x})$. The 4D tensor $\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})$ denotes the $4^{\mathrm{th}}$ central moment of $\bm{t}(\bm{y})$ w.r.t. $p(\bm{y}\,|\,\bm{x})$. These central moments correspond to the cumulants of $\bm{t}(\bm{y})$, i.e. the derivatives of $F$ w.r.t. the natural parameters $\bm{h}(\bm{x})$ of the exponential family. Therefore, the derivatives of $F$ in ^$\hat{\mathcal{I}}_{1}(\theta_{i})$ and ^$\hat{\mathcal{I}}_{2}(\theta_{i})$ can further be written in terms of $\bm{\eta}(\bm{x})$ and $\mathcal{I}(\bm{h}\,|\,\bm{x})$ following the chain rule. Practically, ^$\hat{\mathcal{I}}_{1}$ and ^$\hat{\mathcal{I}}_{2}$ involves computing the Jacobian $\partial\bm{h}(\bm{x})/\partial\theta_{i}$ and the Hessian $\partial^{2}\bm{h}(\bm{x})/\partial^{2}\theta_{i}$, respectively.
In practice, both estimators can be computed via automatic differentiation Paszke et al. (2019); Dangel et al. (2020). In terms of complexity, by restricting to just the diagonal elements $\mathcal{I}(\theta_{i})$, we need to calculate $\mathcal{O}(\dim(\bm{\theta}))$ elements (originally $\mathcal{O}(\dim(\bm{\theta})\times\dim(\bm{\theta}))$ for the full FIM). Although the log-partition function for general exponential family distributions can be complicated, for the ones used in NNs (determined by the loss functions used in optimization) Soen and Sun (2021) the log-partition function $F$ is usually in closed-form; and thus the cumulants $\bm{\eta}(\bm{x})$ and $\mathcal{I}(\bm{h}\,|\,\bm{x})$ can be calculated efficiently.
Indeed, the primary cost of the estimators comes from evaluating the gradient information of the NN, given by $\partial\bm{h}(\bm{x})/\partial\theta_{i}$ and $\partial^{2}\bm{h}(\bm{x})/\partial^{2}\theta_{i}$. The former can be calculated easily. The latter is costly even when restricted to the diagonal elements of the FIM. With the Hessian’s quadratic complexity, in practice approximations are used to reduce the computational overhead Becker et al. (1988); Yao et al. (2020; 2021); Elsayed and Mahmood (2022). In this case, additional error and (potentially) variance may be introduced as a result of the Hessian approximation. Note, the computational cost of the Hessian can still be manageable for the last few layers close to the output. By the chain rule, we only require a sub-computational graph from the output layer to a certain layer to compute the Hessian of that layer. Despite this, there is still a memory cost that scales quadratically with the number of parameters for non-linear activation functions (Dangel et al., 2020).
The high cost of Hessian computation does not justify refraining from using ^$\hat{\mathcal{I}}_{2}$. Depending on the setting (chosen loss function), an estimator’s variance can outweigh the benefits of lower computational costs Soen and Sun (2021). This is especially true when the FIM is used in an offline setting — where the Hessian’s cost can be tolerated — to study, e.g., the singular structure of the neuromanifold Amari et al. (2018); Sun and Nielsen (2019), the curvature of the loss Efron (2018), to quantify model sensitivity Nickl et al. (2023), and to evaluate the quality of the local optimum Karakida et al. (2019; 2021), etc.
@@
To study the quality of ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$ and ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$, it is natural to examine the variance of the estimators Soen and Sun (2021): ${\mathcal{V}}_{j}(\theta_{i}\,|\,\bm{x})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathrm{Var}(\hat{\mathcal{I}}_{j}(\theta_{i}\,|\,\bm{x}))$, where ^$\hat{\mathcal{I}}_{j}(\theta_{i}\,|\,\bm{x})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\left(\hat{\mathcal{I}}_{j}(\bm{\theta}\,|\,\bm{x})\right)_{ii}$ ($j\in\{1,2\}$) is the $i$’th diagonal element of ^$\hat{\mathcal{I}}_{j}(\bm{\theta}\,|\,\bm{x})$. Similar to ^$\hat{\mathcal{I}}_{1}(\theta_{i})$ and ^$\hat{\mathcal{I}}_{2}(\theta_{i})$, ${\mathcal{V}}_{j}(\theta_{i}\,|\,\bm{x})$ and ^$\hat{\mathcal{I}}_{j}(\theta_{i}\,|\,\bm{x})$ depend on the vector $\bm{\theta}$ and are abuses of notation. An estimator with a smaller variance indicates that it is more accurate and more likely to be close to the true FIM. Based on the variance, one can derive sample complexity bounds of the diagonal FIM via Chebyshev’s inequality, see for instance Soen and Sun (2021, Section 3.4).
By its definition, ${\mathcal{V}}_{j}(\theta_{i}\,|\,\bm{x})$ has a simple closed form, which was proved in Soen and Sun (2021) and is restated below.
Lemma 3.1.
$\forall\bm{x}\in\Re^{I}$, $\forall{i}=1,\ldots,\dim(\bm{\theta})$,
$\mathcal{I}(\theta_{i}\,|\,\bm{x})=$ $\;\partial_{i}\bm{h}^{a}(\bm{x})\partial_{i}\bm{h}^{b}(\bm{x})\cdot\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x}),$ (4)
${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})=$ $\;\frac{1}{N}\cdot\partial_{i}\bm{h}^{a}(\bm{x})\partial_{i}\bm{h}^{b}(\bm{x})\partial_{i}\bm{h}^{c}(\bm{x})\partial_{i}\bm{h}^{d}(\bm{x})\cdot\left[\mathcal{K}^{p}_{abcd}(\bm{t}\,|\,\bm{x})-\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})\cdot\mathcal{I}_{cd}(\bm{h}\,|\,\bm{x})\right],$ (5)
${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})=$ $\;\frac{1}{N}\cdot\partial^{2}_{i}{\bm{h}}^{a}(\bm{x})\partial^{2}_{i}{\bm{h}}^{b}(\bm{x})\cdot\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x}).$ (6)
@@
Given a fixed $\bm{x}\in\Re^{I}$, both ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ and ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ have an order of $\mathcal{O}(1/N)$, with $N$ denoting the number of samples of $\bm{y}_{k}$. They further depend on two factors: ① the derivatives of the parameter-output mapping $\bm{\theta}\to\bm{h}$ stored in a $T\times\dim(\bm{\theta})$ matrix, either $\partial_{i}\bm{h}^{a}(\bm{x})$ or $\partial^{2}_{i}{\bm{h}}^{a}(\bm{x})$, where the latter can be expensive to calculate; and ② the central moments of $\bm{t}(\bm{y})$, whose computation only scales with $T$ (the number of output units) and is independent to $\dim(\bm{\theta})$.
@@
From an information geometry Amari (2016) perspective, $\mathcal{I}(\bm{\theta})$, ${\mathcal{V}}_{1}(\bm{\theta})$, and ${\mathcal{V}}_{2}(\bm{\theta})$ are all pullback tensors of different orders. For example, $\mathcal{I}(\bm{\theta})$ is the pullback tensor of $\mathcal{I}(\bm{h})$ and the singular semi-Riemannian metric Sun and Nielsen (2019). They induce the geometric structures of the neuromanifold (parameterized by $\bm{\theta}$) based on the corresponding low dimensional structures of the exponential family (parameterized by $\bm{h}$).
@@
Table 1: Exponential family statistics with eigenvalue upper bounds for moments. For classification, $\sigma(\bm{x})$ denotes the softmax of logit $\bm{h}(\bm{x})$. † denotes exact eigenvalues rather than upper bounds.
@@
Setting Exp. Family Output $\mathcal{Y}$ Sufficient Statistic $\bm{t}(\bm{y})$ UB $\lambda_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))$ UB $\tilde{\lambda}_{\max}(\mathcal{K}(\bm{t}\,|\,\bm{x}))$
Regression (Iso.) Gaussian $\Re^{T}$ ${\bm{y}}$ $1$ $3$
@@ 4 Practical Variance Estimation
To further understand the dependencies of the derivative and central moment terms, the FIM $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ and variances of estimators ^$\hat{\mathcal{I}}_{j}(\theta_{i}\,|\,\bm{x})$ can be bounded to strengthen intuition and to provide a computationally convenient proxy of the interested quantities.
@@
$\frac{1}{N}\cdot\|\partial^{2}_{i}{\bm{h}}(\bm{x})\|^{2}_{2}\cdot\lambda_{\min}(\mathcal{I}(\bm{h}\,|\,\bm{x}))$ $\leq{\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ $\leq\frac{1}{N}\cdot\|\partial^{2}_{i}{\bm{h}}(\bm{x})\|^{2}_{2}\cdot\lambda_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x})),$ (9)
where $\mathcal{M}=\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})-\mathcal{I}({\bm{h}}\,|\,\bm{x})\otimes\mathcal{I}({\bm{h}}\,|\,\bm{x})$; $\lambda_{\min}$ / $\lambda_{\max}$ denotes the minimum / maximum matrix eigenvalue; and $\tilde{\lambda}_{\min},\tilde{\lambda}_{\max}\colon\Re^{T\times T\times T\times T}\rightarrow\Re$ are defined as
$\tilde{\lambda}_{\min}(\mathcal{T})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\inf_{\bm{u}:\|\bm{u}\|_{2}=1}{\bm{u}}^{a}{\bm{u}}^{b}{\bm{u}}^{c}{\bm{u}}^{d}\mathcal{T}_{abcd};\quad\textrm{and}\quad\tilde{\lambda}_{\max}(\mathcal{T})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\sup_{\bm{u}:\|\bm{u}\|_{2}=1}{\bm{u}}^{a}{\bm{u}}^{b}{\bm{u}}^{c}{\bm{u}}^{d}\mathcal{T}_{abcd}.$ (10)
To help ground Theorem 4.1, we summarize different sufficient statistics quantities for common learning settings in Table 1 — with further learning setting implications presented in Section 5. Note that Eqs. 8 and 9 (and many subsequent results) can be further generalized for off-diagonal elements. See Appendix C for details. Compared to prior work Soen and Sun (2021), Theorem 4.1 provides bounds for individual elements of the variance tensors, where the NN weights (the derivatives) and sufficient statistics (the eigenvalues) are neatly disentangled into a product. From a technical point of view, this comes from a difference in proof technique: we utilize variational definitions and computations of eigenvalues to establish bounds whereas Soen and Sun (2021) primarily applies Hölder’s inequality.
We stress that $\tilde{\lambda}_{\min}(\mathcal{T})$ and $\tilde{\lambda}_{\max}(\mathcal{T})$ in Eq. 10 correspond to tensor eigenvalues iff $\mathcal{T}$ is a supersymmetric tensor Lim (2005) (a.k.a. totally symmetric tensor), i.e., indices are permutation invariant. In this case, Eq. 10 is exactly the maximum and minimum Z-eigenvalues. These variational forms mirror the Courant-Fischer min-max theorem for symmetric matrices Tao (2012). In the case of Eq. 8, with $\mathcal{M}=\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})-\mathcal{I}({\bm{h}}\,|\,\bm{x})\otimes\mathcal{I}({\bm{h}}\,|\,\bm{x})$, the tensor is not a supersymmetric tensor in general. Despite this, we note that the lower bound of Eq. 8 is non-trivial. A weaker bound than Eq. 8 can be established based on the Z-eigenvalue of the supersymmetric tensor $\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})$.
@@
The tensor eigenvalue is typically expensive to calculate. However in our case, the eigenvalues $\tilde{\lambda}_{\min}(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x}))$ and $\tilde{\lambda}_{\max}(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x}))$ on the RHS of Eqs. 11 and 12 can be calculated via Kolda and Mayo (2014)’s method with $\mathcal{O}(T^{4}/4!)$ complexity. In this paper, we assume $T$ is reasonably bounded and are mainly concerned with the complexity w.r.t. $\dim(\bm{\theta})$. From this perspective, all our bounds scale linearly w.r.t. $\dim(\bm{\theta})$, and thus can be computed efficiently.
When $\bm{t}(\bm{y})-\bm{\eta}(\bm{x})$ is bounded (e.g. in classification), we can upper bound $\tilde{\lambda}_{\max}\left(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})\right)$ with $\lambda_{\max}\left(\mathcal{I}({\bm{h}}\,|\,\bm{x})\right)$, which is easier to calculate.
@@
Suppose $\|\bm{t}(\bm{y})-\bm{\eta}(\bm{x})\|^{2}_{2}\leq B$. Then, $\tilde{\lambda}_{\max}\left(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})\right)\leq B\lambda_{\max}\left(\mathcal{I}({\bm{h}}\,|\,\bm{x})\right)\leq B^{2}.$
As long as the sufficient statistics $\bm{t}(\bm{y})$ has bounded norm $\|\bm{t}\|_{2}$, we have that $\|\bm{t}(\bm{y})-\bm{\eta}(\bm{x})\|^{2}_{2}\leq 4\|\bm{t}\|_{2}^{2}<\infty$. A similar lower bound can be established for the minimum tensor eigenvalue $\tilde{\lambda}_{\min}\left(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})\right)\geq\lambda^{2}_{\min}\left(\mathcal{I}({\bm{h}}\,|\,\bm{x})\right)$, but this ends up being trivial when applying Corollary 4.2’s lower bound, Eq. 11.
Examining Theorem 4.1 reveals several trade-offs. An immediate observation is that the first order gradients of $\bm{h}(\bm{x})$ correspond to the robustness of $\bm{h}$ to parameter misspecification (w.r.t. an input $\bm{x}$). As such, from the bounds in Eqs. 7 and 8, the scale of $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ and ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ will be large when small shifts in parameter space yield large changes in the output $\bm{h}(\bm{x})$. Another observation is how the spectrum of $\mathcal{I}(\bm{h}\,|\,\bm{x})$ affects the scale of $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ and the estimator variances. In particular, when $\lambda_{\min}\left(\mathcal{I}(\bm{h}\,|\,\bm{x})\right)$ increases, the scale of ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ decreases but the scale of $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ and ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ increases. When $\lambda_{\max}\left(\mathcal{I}(\bm{h}\,|\,\bm{x})\right)$ decreases, then the opposite scaling occurs. With these two observations, there is a tension in how the scale of $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ follows the different variances ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ and ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$. The element-wise FIM $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ follows ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ in terms of the scale of NN derivatives $\|\partial_{i}\bm{h}(\bm{x})\|_{2}$; at the same time, $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ follows ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ in terms of the spectrum of sufficient statistics moment $\mathcal{I}(\bm{h}\,|\,\bm{x})$.